### v1

# On complexity of alternating link equivalence

Abstract.
Link equivalence up to isotopy in a 3-space is the problem that lies at the root of knot theory, and is important in 3-dimensional topology and geometry. We consider its restriction to alternating links, given by two alternating diagrams with $n_{1}$ and $n_{2}$ crossings, and show that this problem has polynomial algorithm in terms of $max\{n_{1},n_{2}\}$. For the proof, we use Tait flyping conjectures, observations from the work of Lackenby, Menasco, Sundberg and Thistlethwaite on alternating links, and algorithmic complexity of some problems from graph theory and topological graph theory.
1. Introduction
One link can be represented by many diagrams, i.e. link projections onto a plane together with information about over and underpasses. It is non-trivial to determine if two link diagrams represent the same link up to isotopy in $\mathbb{R}^{3}$ or $S^{3}$. The following problem lies at the heart of classical knot theory, and was notably mentioned by Turing [42] for one particular link, called the unknot.
link equivalence. Given two link diagrams, do they represent the same link?
Links are important in the study of 3-manifolds, which is reflected, for example, in two classical theorems: Gordon-Luecke [13] and Lickorish-Wallace [26, 27, 43]. In recent years, interest in the algorithmic complexity of problems from 3-dimensional topology and geometry has been growing. Some problems are now known to be decidable or in NP (e.g. [25, 36, 21]), while some problems were proved to be NP-hard ([1, 20, 19, 10, 11, 12, 24]).
In 1960’s and 1970’s, link equivalence was proven to be decidable by Haken and Hemion [16, 15] using theory of normal surfaces. More recent proof, from 2011, is by Coward and Lackenby [8], and uses Reidemeister moves. Such diagrammatic moves were proved to connect any two diagrams of the same link in 1920’s by Alexander, Briggs and Reidemeister [2, 34]. Another proof that link equivalence is decidable can be obtained from the problem of 3-manifold homeomorphism, by passing from a link diagram to the triangulation of the link complement in 3-sphere. As detailed in the work of Kuperberg [21], the 3-manifold homeomorphism problem is known to be decidable due to the proof of Geometrization (conjectured and partly proven by Thurston [41], and completed by Perelman based on the program of Hamilton [32]). While links are not uniquely determined by their complements, one can check whether there is a homeomorphism between the link exteriors taking meridians to meridians (see Corollary 6.1.4 in Matveev’s book [29]), and this is enough for link isotopy. There is also an alternative solution to the homeomorphism problem for hyperbolic link complements by Dahmani and Groves [9], based on work of Sela [37] on isomorphism problem for hyperbolic groups.
Despite such a long and prolific history of study, no lower bounds on complexity of link equivalence are known. Is it NP-hard? Is there a polynomial algorithm? Here, we address its restriction to one infinite family, alternating links, and prove that this restriction is in class P (also called PTIME). Here is the restricted decision problem:
alternating link equivalence. Given two alternating diagrams of a link, do they represent the same alternating link?
Alternating links have been an important class of links: for example, they can be characterized purely topologically [17, 14], and their topological and geometric properties can often be seen or computed directly from a minimal crossing link diagram [30, 22, 39].
Over a century ago, Tait proposed the following conjecture: given prime alternating reduced diagrams $D_{1}$, $D_{2}$ of a link, one can transform $D_{1}$ into $D_{2}$ (up to planar isotopy) by a sequence of diagrammatic moves called flypes. A flype is depicted in Figure 1. The shaded disc labelled $T$ represents a 2-tangle in the terminology of Conway (i.e. a tangle with four free ends). In performing a flype, $T$ is turned by 180 degrees so that the crossing to its left is removed by twisting this tangle, while a new crossing is created to its right. The conjecture was proved in full generality by Menasco and Thistlethwaite as Theorem 1 in [31].
Figure 1. A flype involving a 2-tangle $T$.
We start our discussion of complexity of alternating link equivalence by showing in Section 3 that there are at most $2n^{2}$ flypes between two prime, alternating, reduced and connected diagrams of a link $L$ with $n$ crossings. We then proceed to give an algorithm in Section 4 that compares two such diagrams to determine if they belong to the same link. In Section 5, we check that the number of steps is polynomial in $n$ by reducing certain steps to problems on graphs. This implies that the complexity is polynomial for any two alternating diagrams (without connected, prime, reduced assumptions) due to topological results by Menasco [30].
The fact that alternating link equivalence is polynomial was conjectured, for example, in a recent paper by Samperton [35], as well as in some earlier talks and discussions on knots. The proof is based on known algorithmic complexity of problems from graph theory and topological graphs theory. Other facts that we use come from Tait flyping theorem and Lackenby’s work on volume of alternating links [22] from 2004, though some of these observations were likely known earlier: for example, some appear in the work of Sundberg and Thistlethwaite from 1998 [38].
Note that our algorithm does not immediately apply to alternating links that are not given by alternating diagrams. If a non-alternating diagram is given, there is an algorithm that determines whether the link is alternating given by Howie [17]. If it is, the algorithm produces its alternating diagram up to chirality. The complexity of this algorithm is exponential, since it uses normal surface theory.
It is not yet known if link equivalence in general might be NP-hard. The following related problem was shown to be NP-hard by Koenig and Tsvietkova in [19]:
alternating sublink: Given a diagram of a link $L$ and a positive integer $k$, does $L$ have a $k$-component sublink that is alternating?
This can be compared with the fact that subgraph isomorphism is NP-hard [5].
2. Preliminaries: alternating links, their diagrams, and graphs
2.1. Some basics from knot theory.
Let $L$ be an alternating link, i.e. such that its minimal crossing diagram is alternating. In an alternating diagram, overpasses and underpasses alternate as one travels along the link. Recall that a link diagram is prime if any simple closed curve that intersects the diagram in exactly two points has only an unknotted arc on one of its sides. A diagram is reduced if it does not contain nugatory (i.e. reducible) crossings. Consider a connected prime alternating reduced diagram $D$ of $L$ with $n$ crossings. Due to Menasco’s results [30], the link $L$ must be topologically prime and non-split. Reduced alternating diagram must also have the minimal crossing number, say $n$, due to another Tait conjecture, proved by Kauffman, Murasugi, and Thistlethwaite [18, 33, 40].
An edge of a link diagram is its segment (an arc) from a crossing to the nearest crossing. Any link diagram with $n$ crossings has $2n$ edges. A region of a link diagram (further diagram region) is bounded by crossings and edges, and has none in its interior, i.e. is complementary. A diagram region bounded by only two edges and two crossings is a bigon. Any maximal sequence of bigons in $D$, where each two consecutive bigons share a crossing, is called a twist. For a twist, a subset of consecutive bigons is a subtwist. A subtwist can be maximal as well, i.e. equal to a twist.
2.2. Squares, tangles, and flypes.
Let us recall the terminology from Section 4 of Lackenby’s paper [22]. A square is a simple closed curve that intersects a link diagram four times. If the intersections lie on four different edges of the link diagram, the square is essential. A square is characteristic if it is essential, does not separate off a single crossing on either side, and any other square can be isotoped off it, where the isotopy is planar (and hence one cannot isotop through the crossings of $D$). A characteristic collection of squares is defined as follows: take one isotopy class of each characteristic square, and isotop them so that they are all disjoint.
An essential square corresponds to a 2-tangle that has at least one crossing. Such a tangle can be the tangle $T$ in the definition of a flype. If $T$ has more than one crossing, we will say that the flype is non-trivial. A trivial flype (for a tangle $T$ with 0 or 1 crossings) does not change an alternating link diagram. Hence in what follows, we will only need to consider non-trivial flypes.
For a collection $\mathcal{C}$ of disjoint essential squares, a product region is defined as follows. Color the essential squares and their interior in blue. Consider $D$ without the blue interior. The regions bounded by arcs of $D$ (edges) and blue arcs on the boundary of squares can be colored in black and white in checkerboard fashion, since diagram regions of any reduced alternating diagram can be colored this way. If for color X (X is black or white), all regions of color X intersect crossings of $D$ and blue squares in only two connected components (e.g. in two crossings, or in a crossing and a blue arc), call the union of all regions of color X a complementary product region for $\mathcal{C}$. Lackenby demonstrated in Lemma 8 of [22] that a product region consists of twists and tangles beaded on two strands of a link as in Figure 2. This construction was also considered earlier by Sundberg and Thistlethwaite [38], and called a flyping circuit there.
Figure 2. Essential squares (in blue) and their complementary regions. In this figure, one can see a product region.
A flype on any essential square of $\mathcal{C}$ will both flip the square and move it one crossing to the left or right along the product region. A sequence of flypes can exchange such a square with a twist, or a subtwist in the same product region, but cannot exchange the squares between themselves.
We will use the following lemma.
Lemma 1.
[22, Lemma 9 and Lemma 10] Let $\mathcal{C}$ be a collection of disjoint non-parallel essential squares. It is a characteristic collection if and only if all of the following hold:
  1. (1)
    any essential square in the complement of $\mathcal{C}$ lies in a product complementary region, or is parallel to a square of $\mathcal{C}$, or encloses a single crossing;
  2. (2)
    if two product complementary regions of $\mathcal{C}$ are adjacent, they have incompatible product structures;
  3. (3)
    no square of $\mathcal{C}$ encloses a single crossing.
The condition (2) in the above lemma means the following situation does not occur: a square of a characteristic collection has two complementary regions on different sides of it (inside and outside), both of which are product regions, and their union is again a product region.
2.3. Graphs and plane graphs.
We recall some classical problems about graphs.
Suppose we have a graph $\Gamma=(V,E)$ and function $f:V\rightarrow A$, where $A$ is a section of $\mathbb{Z_{+}}$, i.e. the initial segment of $\{1,2,3,...\}$. We then call $f$ a coloring function, $\Gamma$ a vertex colored graph, and $f^{-1}(i)$ the color class of $i$ or of a vertex colored with $i$. We say that color valence of $\Gamma$ is less or equal than $d$ with respect to $f$, if for every $v$ and $i$, either the number of neighbors or the number of non-neighbors of $v$ in $f^{-1}(i)$ is less or equal than $d$.
The following is a well-studied decision problem.
vertex colored graph isomorphism with bounded color valence. Given two vertex colored graphs with bounded color valence, are they isomorphic?
The isomorphism problem for colored grahs with bounded color valence is known to be solvable in polynomial time. For example, Prop. 4.5 of [3] assures that canonical form of a vertex colored graph with color valence at most $d$ can be computed in $|V|^{P(d)}$ steps, where $P(d)$ is a polynomial that depends only on $d$. The graph canonization problem is at least as computationally hard as the graph isomorphism problem, so vertex colored graph isomorphism with bounded color valence has computational complexity at most $|V|^{P(d)}$ as well. This can be compared with graph isomorphism, which is not known to lie in P (and thus the problem for colored graphs is potentially easier).
A pseudograph is a graph that allows multiple edges between two vertices as well as loops, i.e. edges that connect a vertex to itself. pseudograph isomorphism with bounded valence can be reduced to graph isomorphism with bounded valence by adding extra vertices subdividing some edges. The resulting graph is called a subdivision graph. If we start with a pseudograph with $n$ vertices and $2n$ edges, no loops, some double edges, and each vertex having valence 4, the subdivision graph is obtained in polynomial number of steps in terms of $n$, has linearly many edges and vertices in terms of $n$, and valence (and hence color valence) bounded above by 4. In particular, graph isomorphism with bounded valence is at least as hard as pseudograph isomorphism with bounded valence, and the former problem is in P as was proved by Luks [28]. Similarly, vertex colored graph isomorphism with bounded color valence is at least as hard as vertex colored pseudograph isomorphism with bounded color valence, and so both are in P.
Recall that a graph can be embedded in the plane if it can be drawn in the plane without crossings between edges. Such graphs are called planar. There can be many plane embeddings of one planar graph. In some literature, a graph together with a fixed embedding is called a plane graph. Such graphs are considered up to isotopy, which is a stronger notion than graph isomorphism: in particular, such isotopy is the map between the graphs that can be extended to a map from the plane to itself, isotopic to the identity. We will call the problem of comparing such graphs plane graph isotopy: it has been studied extensively in topological graph theory, and is known to be solvable in polynomial time. (For more on this problem, see, for example, [7, 6, 4]) .
Using the known algorithms, one obtains that comparing two plane vertex colored plane graphs or pseudographs up to isotopy preserving colors is in class P. For example, one way to do it for vertex colored connected plane graphs $\Gamma_{1}$ and $\Gamma_{2}$ is as follows.
Pick an arbitrary vertex $v_{1}$ in $\Gamma_{1}$ and an edge $e_{1}$ adjacent to it. For each vertex $v_{2}$ in $\Gamma_{2}$, with the color $f(v_{1})=f(v_{2})$, and an incident edge $e_{2}$, consider a map from the pair $(v_{1},e_{1})$ to the pair $(v_{2},e_{2})$. For two planar embeddings to be isotopic, the cyclic order of the edges around equivalent vertices should be the same. Hence, by mapping a vertex and an edge, we uniquely (up to the change of clockwise/anticlockwise direction) determine how the edges adjacent to $v_{1}$ are mapped into the edges adjacent to $v_{2}$. Once the map has been extended to all edges adjacent to $v_{1},v_{2}$, this also determines the map for vertices on the other end of the edges that were mapped. We first compare colors for these newly mapped vertices, and, if they coincide, proceed. Suppose such a vertex $v_{1}^{\prime}$ in $\Gamma_{1}$ is mapped to $v_{2}^{\prime}$ in $\Gamma_{2}$. Since for $v_{1}^{\prime},v_{2}^{\prime}$ at least one incident edge has already been mapped, we can use it to map other edges respecting the cyclic order again. Proceed by extending the map locally this way to all vertices and edges. This process has at most linearly many steps in terms of $|V|$. If for one choice of $(v_{2},e_{2})$, we were able to extend the map to the whole graph, there is an isotopy of plane graphs preserving colors. Otherwise, after trying all the possibles choices for $(v_{2},e_{2})$, there is no such isotopy. The total number of steps is at most $O(|V|^{2})$.
3. Bound on the number of flypes
Bounds on purely diagrammatic moves that take one link diagram to another one have been a rich topic of study. Previous work concerns mostly Reidemeister moves (e.g. [8, 23]), but it is perhaps useful to have a simple bound on the number of flypes as well. More importantly for us, the facts needed for the polynomial algorithm are established in the proof of the following proposition.
Proposition 2.
Suppose we are given two diagrams of a link $L$, with each diagram being connected, prime, alternating, and reduced. Then at most $2n^{2}$ flypes are needed to transform one diagram to another, where $n$ is the crossing number.
Proof.
Denote the two diagrams given in the theorem by $D$ and $D^{\prime}$. As noted above, by one of Tait flyping conjectures (proved in [18, 33, 40]), the crossing number of these diagrams is the same. Consider a characteristic collection $\mathcal{C}$ for $D$.
For every tangle $T$ that can be flyped, there is an essential square that encircles it. From Lemma 1 (1), any essential square that is not in $\mathcal{C}$ up to planar isotopy either encircles a single crossing or lies in a complementary product region. If we want to count all possible flypes, we are not interested in any squares that encircle a single crossing (call them trivial squares), since they give rise to trivial flypes. A square that lies in a product region encircles a subset of squares of $\mathcal{C}$ and subtwists of the product region. Hence in $D$, we can see all possible tangles that we can flype (i.e. all potential tangles $T$ in the definition of a flype): such a tangle is either encircled by a square of $\mathcal{C}$ or by a non-trivial essential square in a product region. Moving subtwists along a product region either does not change the diagram, or is equivalent to moving some squares from the characteristic collection. Hence we may ignore non-trivial essential squares that do not encircle any squares of $\mathcal{C}$ in the product region. Therefore, we may assume that in the definition of a flype, the tangle $T$ consists of squares $c_{1},...,c_{k}$, with $k\geq 1$, of $\mathcal{C}$, that share a product region $P$, and possibly some subtwists from $P$.
If $k=1$ and there are no subtwists from $P$ in $T$, call a flype that involves the tangle $T$ simple. Otherwise call it a concatenation flype. A concatenation flype can be substituted by a sequence of simple flypes. Hence, in what follows, we may consider only simple flypes, and count them for an upper bound. Suppose we have a sequence of diagrams $D=D_{1},D_{2},D_{3},...,D_{m}=D^{\prime}$, where $D_{i+1}$ is obtained from $D_{i}$ by a simple flype. We need to bound $m$.
Once a flype is applied to a diagram, some squares of $\mathcal{C}$ might move with the flype along their product regions, but otherwise the squares do not change, and a characteristic collection is still a characteristic collection after a flype, in the resulting diagram. The proof of this is straightforward and can be found in the first paragraph of Lemma 4 in [22]. In particular, $|\mathcal{C}|=|\mathcal{C}^{\prime}|$.
Now let’s count possible flypes that take us from a diagram $D$ to a diagram $D^{\prime}$ of $L$. In the initial diagram $D$, we have up to $n$ squares of $\mathcal{C}$. Indeed, each square $c$ of $\mathcal{C}$ has more than one crossing, but $c$ might have some other squares of $\mathcal{C}$ inside it. So we assume that for a square $c$, there is at least one crossing $x$ such that $x$ is not contained in other squares of $\mathcal{C}$ that do not contain all of $c$. There are also up to $n$ product regions in $D$, since each such region should contain at least one essential square or crossing.
A simple flype can move one of $n$ squares of $\mathcal{C}$, say $c$, along its product region. Note that $c$ might have one product region, say $P_{1}$, on one side of $c$ (say, inside the square $c$), and another product region, say $P_{2}$, on the other side of $c$ (say, outside of $c$). Each product region in a link diagram has up to $n$ crossings outside of its squares, and hence $c$ can be moved up to $n$ times by a simple flype. Hence, when we choose square $c$ of $\mathcal{C}$, we have $n$ options; when we choose inside/outside product region for $c$, we have 2 options ($P_{1},P_{2}$); and we can move $c$ at most $n$ times in each product region. This yields $2n^{2}$ possible simple flypes, each potentially taking us to one of the diagrams in the sequence $D=D_{1},D_{2},D_{3},...,D_{m}=D^{\prime}$.
If there are more flypes, they must reverse what has been done by previous flypes, since we ran out of new options. Hence $m\leq 2n^{2}$. ∎
As corollary of the theorem, the sequence of flypes can serve as a polynomial certificate, putting alternating link equivalence in complexity class NP. Note that to specify a flype, it is enough to specify four edges for the square that encircles the tangle $T$, and specify a crossing outside of $T$. So to specify a sequence $P(n)=2n^{2}$ flypes, a sequence of $4P(n)$ edges and $P(n)$ crossings is needed. The NP upper bound on complexity will be improved to P in the next sections.
4. Algorithm for alternating link equivalence
In this section, we give a simple algorithm for alternating link equivalence. In this section, we assume that alternating link diagrams $D_{1},D_{2}$ are prime, reduced, and connected. Once we have an algorithm for such diagrams, the algorithm for any two alternating diagrams is discussed, for example, in [35], based on the work of Menasco [30].
4.1. Basis for the algorithm
As noted above, the crossing numbers of $D_{1}$ and $D_{2}$ must be equal if the links are isotopic.
The following facts follow from the definition of a flype or are explained in Proposition 2 proof.
  1. (1)
    Characteristic collection $\mathcal{C}_{i}$ of $D_{i}$ is invariant under flypes: its squares may move inside their product regions, but no squares are removed or added to $\mathcal{C}_{i}$ when a flype is performed.
  2. (2)
    The 2-tangles (i.e. their projections in a link diagram) inside the squares of $\mathcal{C}_{i}$ do not change beyond planar isotopy when flypes are performed.
  3. (3)
    Every non-trivial flype can be substituted by a finite sequence of simple flypes such that for each simple flype, there is a square of $\mathcal{C}_{i}$ encircling the tangle $T$ that will be flyped.
  4. (4)
    Flypes cannot exchange squares of $\mathcal{C}_{i}$ in a product region.
  5. (5)
    A product region $P$ is uniquely determined by tangles inside the squares of $\mathcal{C}_{i}$ that $P$ contains, the cyclic order and number of squares (this number is called weight in [38]), and by the number of crossings in $P$ outside of the squares of $\mathcal{C}_{i}$ (call them free crossings of $P$).
Note that two equivalent product regions might look different, since we do not pay attention to the number of crossings between each two tangles, and only compare the total weight of the regions. But this is easily adjusted by flypes.
Therefore, to compare $D_{1}$ with $D_{2}$, it is enough to compare the tangles or product regions encircled by the squares of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, and then to compare how these squares “fit together” to make a link diagram. They might “fit together” in either product region or some other structure. (While we will not be using this, such structures are considered in [38] under the name “templates”, with pictures and examples.)
Squares of characteristic collection $\mathcal{C}_{i}$ may encircle tangles that have other squares of $\mathcal{C}_{i}$ in them, non-parallel to the initial square. If a link diagram is on a plane, each square has two sides: inside and outside. We call a side of a square of $\mathcal{C}_{i}$ inner if there are no other squares of $\mathcal{C}_{i}$ in the tangle that it encircles on that side. If there is only one square in a characteristic collection, it has two inner sides. We start by comparing tangles on the inner sides of squares, coloring them, then comparing tangles and regions which are right outside of them and coloring them next, etc., until we finally obtain two colored pseudographs of bounded valence instead of the two link diagrams we started with.
4.2. Algorithm
Input: two connected prime alternating reduced diagrams $D_{1},D_{2}$ with crossing numbers $n_{1},n_{2}$.
Step 1. If $n_{1}\neq n_{2}$, stop, the links are not isotopic. If $n_{1}=n_{2}$, go to step 2.
Step 2. Detect a characteristic collection of squares $\mathcal{C}_{i}$ in $D_{i},i=1,2$
Step 3. Determine all inner sides of squares of $\mathcal{C}_{i},i=1,2$, and label them with $i_{1}$ for “inner”. Then label the squares that contain the already labelled squares but not other squares by $i_{2}$ on the side where the $i_{i}$ squares are. Continue until all squares are labelled this way. (Note: we might have a square labelled by $i_{1}$ on both sides, and then it is the only square in $D_{i}$.) See Figure 3 (A) for an example of such labelling.
Step 4a. When $j=i_{1}$, to compare tangles $T_{1},T_{2}$, make a pseudograph $\Gamma_{i}$ out of $T_{i},i=1,2,$ as follows. Ignore the parts of $D_{i}$ outside of $T_{i}$. In $T_{i}$, put a vertex instead of every crossing, and an edge instead of every diagram edge. For the four free ends of $T_{i}$, put a vertex of valence 1 on each (in $D_{i}$, the vertex lies on the intersection of the square encircling $T_{i}$ and the diagram edge), and give them labels $N,S,E,W$ in $T_{1}$. Then compare the resulting plane pseudographs of valence at most 4. If two pseudographs are the same, compare crossing information for the respective tangles (since the tangles are alternating, it is enough to compare just one pair of crossings, one crossing from one tangle, and one from another). If it is the same, the tangles have the same color. Label the four vertices on tangle $T_{1}$ “free ends” by $N,S,E,W$ according to the equivalence with the vertices of tangle $T_{1}$.
We now have all tangles on the $i_{1}$-sides of squares colored according to the equivalence of tangles. Make each colored tangle together with the square that encircles it into a colored graph vertex, otherwise leaving the diagram $D_{i},i=1,2$, untouched. Denote the resulting graph/diagram by $D^{\prime}_{i},i=1,2,$ respectively. The vertices of $D^{\prime}_{i},i=1,2,$ that are not colored all get the same (not previously used) color.
Step 4b. Repeat the following until we exhaust all values from $i_{2},...,i_{k}$ for $j$. Change $j$ to the next label in the list of $i_{2},....i_{k}$. By now, there are colored vertices inside any square, say $c$, that we consider. Then $c$ must encircle a product region by Lemma 1 (and all of it, since the product regions inside and outside of $c$ are not compatible by the lemma). We compare two product regions $P_{1},P_{2}$ in $D_{1}^{\prime},D_{2}^{\prime}$ respectively, as follows. Compare the list of colors of tangles inside $P_{1},P_{2}$ up to (possibly reverse) cyclic order, and compare the total number of free crossings for $P_{1},P_{2}$. If these numbers and lists are equal, also compare the lists of $N,S,W,E$ labels for free ends of equivalent tangles inside $P_{1},P_{2}$, and the crossing information for free crossings. If all coincides, the product regions are considered equivalent. Compare all pairs of product regions for the label $j$, and give the equivalent ones the same color. Make these newly colored product regions into new colored vertices, changing $D^{\prime}_{i},i=1,2$. Label the four vertices of $P_{1},P_{2}$ “free ends” by $N,S,E,W$ according to the equivalence between $P_{1}$ and $P_{2}$.
Figure 3. In Figure (a), there is an example of a link diagram $D_{1}$ (in black) with a characteristic collection $\mathcal{C}_{1}$ of squares (in blue), where inner sides of squares are labeled by $i_{1},i_{2}$. Substitute 2-tangles on the sides of squares of $\mathcal{C}_{1}$ labeled by $i_{1}$ with colored graph vertices (red and green). Substitute remaining crossings of 2-tangles on the sides of squares of $\mathcal{C}_{1}$ labeled by $i_{2}$ with black vertices.
Step 5. At this step, we have no more squares. Make a colored pseudograph of what remains of $D^{\prime}_{i},i=1,2$, as in Step 4a. Compare two such vertex colored plane pseudographs.
Making pseudographs is illustrated in an example in Figure 3 (B) for $j=i_{2}$, with two blue squares encircling colored pseudographs made out of tangles.
5. Algorithmic complexity
Theorem 3.
The problem alternating link equivalence is in class P. In particular, the above algorithm has a polynomial number of steps in terms of $\text{max}\{n_{1},n_{2}\}$.
Proof.
The algorithm in Section 4 is for connected prime alternating reduced diagrams. We first discuss its complexity, and then its generalization to any alternating diagrams.
Complexities of steps 1-5 need to be added. If $n_{1}\neq n_{2}$, the algorithm stops at step 1. So we may assume that $n_{1}=n_{2}=n$.
Step 1 is comparison of two natural numbers.
Step 2 can be subdivided in two substeps, each with polynomial complexity, and their complexities need to be added.
First, detect all essential squares. For this, take all edges of $D_{i},i=1,2$. There are $2n$ of them for each diagram. Then take all (unordered) combinations of 4 edges out of them: there are (${2n\choose 4}=O(n^{4})$ of them. For each quartet of edges, check if the curve intersecting them is connected. If it is, it is a square. Check if there is more than one crossing inside and outside this square. If yes to both, then the square essential. Therefore, all essential squares can be detected in polynomial number of steps in terms of $n$, and there are at most $O(n^{4})$ of them.
Second, choose characteristic squares out of the essential squares. For this, for each pair of essential squares (there are at most $O(n^{8})$ such pairs) check if they intersect as curves, and if the intersection contains a crossing. If the answer for a pair is “yes”, the squares are not characteristic. If the answers for a square have been “no” in every pair, then it is a characteristic square. All together, this takes polynomial number of steps in terms of $n$.
Note that there are at most $n$ characteristic squares. Indeed, such a square $c$ must contain at least one crossing that is not contained in other characteristic squares that do not fully contain $c$.
In Step 3, for each square of the characteristic collections $\mathcal{C}_{i},i=1,2$, check if there is another square fully inside it: e.g., by checking if a square shares all crossings with another square. For each of at most $4n^{2}$ pairs of sides of squares, this requires checking at most $n$ crossings, and hence yields at most $4n^{3}$ steps. This way we detect and label all $i_{1}$ sides of squares of $\mathcal{C}_{i},i=1,2$. Then for each $i_{1}$ square, consider its non-labelled side, and check whether it shares crossings with any other non-labeled side of another square. Again there are up to $4n^{2}$ pairs of sides of squares, and up to $n$ crossings for each pair to check. We have at most $4n^{3}+4n^{3}$ steps to detect $i_{1}$ and $i_{2}$ squares. Continue at most $n$ times, get at most $n$ by $4n^{3}$ steps to label all sides of squares by $i_{1},i_{2},i_{3},....i_{k},k\leq n$.
Step 4 is repeated until every two tangles or product regions encircled by the squares of $\mathcal{C}$ are compared, where a tangle or a product region can lie on one of two sides of a square. So we have at most $2n$ times $2n$ pairs of tangles or product regions to compare. When we create pseudographs from fragments of a link diagram, we do this in polynomial number of steps. In the resulting pseudographs, there are no loops and there are at most two double edges between two vertices. Due to properties of a link diagram, there are at most $n$ pseudograph vertices (or $2n$ for a subdivision graph), the valence of each vertex is at most $4$, there are up to $2n$ edges for pseudograph (or $3n$ for a subdivision graph). The total number of colors is at most $n$, since we have at most $n-1$ distinct tangles, and can add one more color that is used for any pseudograph vertex that has not been obtained from a colored tangle.
In Step 4a, for $j=i_{1}$, the tangle comparison is polynomially reduced to plane pseudograph isotopy with bounded valence, which has polynomial complexity as explained in subsection 2.3.
In Step 4b, for $j=i_{2},...,i_{k}$, the comparison for each pair of product regions is reduced to comparison of two ordered lists of colors, comparison of two numbers (i.e. the weights of product regions), and comparing $N,S,W,E$ labels for each of at most $n$ equivalent pairs of equivalent tangles. This can be done in polynomially many steps.
The polynomials for these two substeps must be multiplied by $4n^{2}$, i.e. the upper bound on number of pairs of tangles or product regions to compare, and added.
Step 5 is polynomially reduced to vertex colored plane pseudograph isotopy with bounded color valence. It is known to have polynomial complexity as explained in subsection 2.3.
Now once there is a polynomial algorithm to compare connected prime alternating reduced diagrams, one can compare any two alternating diagrams in polynomial time. This is based on the following facts: (i) if an alternating diagram is not reduced, one can remove nugatory crossings by a 180-degree twist in ambient 3-space, and there are at most $n$ nugatory crossings; (ii) Menasco [30] proved that connected alternating diagrams represent non-split links; (iii) Menasco also proved that prime connected alternating diagrams represent prime links. Subsection 2.1 of [35] carefully describes how to use this to pass from any alternating diagrams to connected prime alternating reduced diagrams in polynomial number of steps.
6. Acknowledgements
We thank March Lackenby for helpful discussions, and Arnaud De Mesmay for helpful correspondence and encourangement. Tsvietkova is supported by NSF DMS-2005496 and NSF CAREER DMS-2142487.